A source of sound producing wavelength 50cm is moving away from a stationary observer with 1/5th the speed of sound. Then the wavelength of sound as measured by observer is

Grade:12

1 Answers

User
6 years ago
Wavelength =50 cmVelocity =vSource velocity =v/5Apparent wavelength =[50*(v-v/5)/v] Apparent wavelength =40 cm